### v1

# PO-CKAN:Physics Informed Deep Operator Kolmogorov Arnold Networks with Chunk Rational Structure

Abstract
We propose PO-CKAN, a physics-informed deep operator framework based on Chunkwise Rational Kolmogorov–Arnold Networks (KANs), for approximating the solution operators of partial differential equations. This framework leverages a Deep Operator Network (DeepONet) architecture that incorporates Chunkwise Rational Kolmogorov–Arnold Network (CKAN) sub-networks for enhanced function approximation. The principles of Physics-Informed Neural Networks (PINNs) are integrated into the operator learning framework to enforce physical consistency. This design enables the efficient learning of physically consistent spatio-temporal solution operators and allows for rapid prediction for parametric time-dependent PDEs with varying inputs (e.g., parameters, initial/boundary conditions) after training. Validated on challenging benchmark problems, PO-CKAN demonstrates accurate operator learning with results closely matching high-fidelity solutions. PO-CKAN adopts a DeepONet-style branch–trunk architecture with its sub-networks instantiated as rational KAN modules, and enforces physical consistency via a PDE residual (PINN-style) loss. On Burgers’ equation with $\nu=0.01$, PO-CKAN reduces the mean relative $L^{2}$ error by approximately 48% compared to PI-DeepONet, and achieves competitive accuracy on the Eikonal and diffusion–reaction benchmarks.
keywords:
Physics Informed Neural Network (PINN) , Deep Operator network (DeepONet) , Kolmogorov–Arnold Networks (KANs) , Rational activation functions , Partial Differential Equations
journal: Journal of Computational Physics
[…]
1 Introduction
The Physics-Informed DeepONet (PI-DeepONet) [29] is an emerging machine learning framework designed to efficiently solve entire families of parameterized partial differential equations (PDEs) [35, 36, 20, 37, 42, 26]. Unlike traditional Physics-Informed Neural Networks (PINNs) [43], which solve a single problem, PI-DeepONet is designed to approximate the underlying solution operator. This operator constitutes a universal mapping from any given input function (e.g., boundary conditions, forcing terms) to its corresponding solution function [40, 13, 7, 43, 48, 58, 24, 47]. It achieves this through a unique ”Branch–Trunk” architecture and ensures that its learning process adheres to fundamental physical laws by incorporating the governing equation residuals into its loss function. Its most prominent advantage is its ”train once, predict many” efficiency, which drastically reduces computational expense in scenarios requiring large-scale repeated solutions. A promising frontier in this area involves enhancing the PI-DeepONet framework by replacing its MLP components with a more expressive architecture, an approach that has already shown success [12, 29, 34].
Kolmogorov-Arnold Networks (KANs) [32, 25, 46, 31] present a compelling alternative to conventional Multilayer Perceptrons (MLPs) [23] and are thus well-suited for this application. KANs, which are theoretically grounded in the Kolmogorov-Arnold Representation Theorem [39], introduce a significant architectural departure from MLPs. Unlike MLPs, which pair fixed activation functions on the nodes with learnable weights and biases [5, 50], KANs place learnable activation functions, originally parameterized as B-splines [52, 19], on the edges of the network. This fundamental difference in architecture yields notable benefits, such as improved model interpretability and more robust performance on tasks characterized by noisy data or requiring continual learning [56, 11, 18, 30, 53, 22, 41].
Importantly, this theoretical power has been validated in the practical application of solving differential equations. For instance, KANs have been successfully integrated into PINNs to solve the Poisson equation [57], and a KAN-based operator network, DeepOKAN [2], was introduced to simulate sinusoidal wave solvers and address orthotropic elasticity problems using Radial Basis Functions (RBFs). Both studies confirmed that KANs significantly outperform traditional MLP architectures. However, despite this demonstrated superiority, a critical flaw obstructs their widespread adoption. The total number of learnable parameters in KANs grows quadratically with network width. This severe scalability bottleneck renders a naive integration into the DeepONet framework computationally impractical for most operator learning tasks [14, 8, 15], creating a significant gap between the architecture’s theoretical potential and its practical application.
To overcome this fundamental challenge, we introduce our primary technical innovation: the Chunkwise Rational KAN (CKAN). This architecture is specifically designed for parameter efficiency. It employs a chunk-wise parameter sharing mechanism, where chunks of connections share a common activation function, while individual edges retain unique scalar weights. To further enhance performance, we leverage the power of rational functions for these activations. Foundational work has established that rational functions, which have the form $R(x)=P(x)/Q(x)$, are highly effective approximators, particularly for functions with complex features such as singularities or steep gradients [6, 10, 49, 27]. Integrating such learnable rational functions into the KAN structure presents a compelling direction for creating more powerful and efficient neural networks [3, 38]. Our CKAN design actualizes this concept. This novel combination of chunking and rational activations breaks the quadratic scaling bottleneck, making the expressive power of KANs computationally feasible for large-scale operator learning.
By integrating our efficient CKAN architecture into the physics-informed DeepONet structure, we construct our final model: the Physics-Informed Deep Operator Chunk-rational KAN (PO-CKAN). This framework inherits the operator learning paradigm from its DeepONet foundation while leveraging the enhanced and scalable expressiveness of our CKAN architecture to achieve superior accuracy and greater computational efficiency. The primary contributions of this study are:
  • 1.
    We introduce an efficient rational activation function, a lightweight alternative to splines that enhances computational efficiency while maintaining high expressive power.
  • 2.
    We design the Chunkwise Kolmogorov-Arnold Network (CKAN), a novel architecture that employs a chunk-wise parameter sharing mechanism to break the quadratic ($O(N^{2})$) scaling bottleneck of traditional KANs.
  • 3.
    We propose the Physics-Informed Operator CKAN (PO-CKAN) by integrating our scalable CKAN into the DeepONet architecture, which to our knowledge marks the first successful incorporation of KAN-like expressivity into physics-informed operator learning in a scalable manner.
  • 4.
    We demonstrate the effectiveness and versatility of the PO-CKAN framework across a comprehensive suite of four benchmarks. Our model successfully tackles diverse problem classes—from time-evolution systems (e.g., Burgers’ equation) to boundary-value problems (e.g., Eikonal equation)—and consistently achieves a superior accuracy-efficiency trade-off compared to traditional DeepONet baselines.
The structure of this paper is as follows. Section 2 introduces our proposed framework, the Physics-Informed Deep Operator Chunk-rational KAN (PO-CKAN). Subsequently, Section 3 validates the framework’s performance through a series of numerical experiments on benchmark problems, including the Eikonal Equation, Burgers’ Equation, a Fractional PDE, and a Diffusion-reaction System. Finally, Section 4 discusses our conclusions and outlines potential directions for future research.
2 Methodology
2.1 Preliminaries
2.1.1 Deep Operator Network (DeepONet)
The DeepONet framework is designed to approximate solution operators for PDEs. It employs a distinctive “Branch–Trunk” architecture to learn the mapping from an input function, $u(y)$, to a solution function, $s(x)$. The core components are:
  • 1.
    Branch Network: Encodes the input function $u(y)$ into a $p$-dimensional feature representation, $b=[b_{1},\dots,b_{p}]\in\mathbb{R}^{p}$.
  • 2.
    Trunk Network: Processes the spatio-temporal coordinates, $x$, to learn a set of $p$ coordinate-dependent basis functions, $t(x)=[t_{1}(x),\dots,t_{p}(x)]$.
The final solution, $s(x)$, is then reconstructed as the dot product between the outputs of the two networks:
$G(u)(x)\approx\sum_{k=1}^{p}b_{k}\cdot t_{k}(x)=\langle b,t(x)\rangle$ (1)
2.1.2 Kolmogorov–Arnold Network (KAN)
Building upon the concepts introduced earlier, we now formally define the Kolmogorov–Arnold Network (KAN) architecture and its mathematical underpinnings.
A foundational result in function approximation theory, the Kolmogorov–Arnold representation theorem, offers a powerful alternative to the conventional paradigms of neural network design [21]. The theorem reveals that the apparent complexity of multivariate continuous functions is reducible; that is, any such function on a bounded domain can be expressed as a finite composition of single-variable functions linked by addition. For a specific smooth function $f:[0,1]^{n}\to\mathbb{R}$, the theorem guarantees that the following representation holds:
$f(\mathbf{x})=f(x_{1},\dots,x_{n})=\sum_{q=1}^{2n+1}\Phi_{q}\left(\sum_{p=1}^{n}\phi_{q,p}(x_{p})\right).$ (2)
In this formulation, both the inner functions $\phi_{q,p}:[0,1]\to\mathbb{R}$ and the outer functions $\Phi_{q}:\mathbb{R}\to\mathbb{R}$ are univariate. This remarkable theoretical statement suggests that simple addition is the only multivariate operation required. However, while theoretically profound, the functions guaranteed by the original theorem can be non-smooth and pathological, which has historically limited its direct application and hindered attempts to develop practical, learnable models based on it.
The Kolmogorov–Arnold Network (KAN) [32] provides a learnable framework for the theorem by modeling its univariate functions with B-splines [17, 45, 44, 1, 9]. Fundamentally differing from MLPs, KANs employ learnable activation functions on the edges of the network rather than fixed activations on the nodes.
A KAN is constructed as a sequence of layers defined by its shape $n_{0},n_{1},\dots,n_{L}]$. A matrix of learnable activation functions connects each layer to the next. The output of a neuron in a given layer is the sum of the outputs from all of its incoming functions. This process is described by the equation:
$x_{l+1,j}=\sum_{i=1}^{n_{l}}\phi_{l,j,i}(x_{l,i}),\quad j=1,\dots,n_{l+1}.$
The full network is therefore a composition of these function matrices, $\mathbf{\Phi}_{l}$:
$\text{KAN}(\mathbf{x})=(\mathbf{\Phi}_{L-1}\circ\mathbf{\Phi}_{L-2}\circ\cdots\circ\mathbf{\Phi}_{0})(\mathbf{x}_{0}).$
This structure is fundamentally different from a Multilayer Perceptron (MLP). An MLP alternates between linear transformations, $\mathbf{W}$, and a single, fixed non-linear function, $\sigma$:
$\text{MLP}(\mathbf{x})=(\mathbf{W}_{L-1}\circ\sigma\circ\mathbf{W}_{L-2}\circ\sigma\circ\cdots\circ\mathbf{W}_{0})(\mathbf{x}).$
The key distinction, therefore, is that KANs embed both linear transformations and nonlinearities into the learnable function matrices $\mathbf{\Phi}$.
This design gives KANs superior expressive power and interpretability. However, it also creates a major challenge. The activation functions are typically parameterized using B-splines. This B-spline parameterization yields a large number of parameters and incurs a substantial computational cost. Consequently, the standard KAN design does not scale well. A more efficient and parsimonious architecture is therefore needed for large-scale applications.
2.2 The Proposed PO-CKAN Framework
This section details the architecture and training objective of the Physics-Informed Deep Operator Chunk-rational KAN (PO-CKAN). We begin by describing the overall structure before providing a mathematical formulation of the novel CKAN layer. Finally, we define the components of the physics-informed loss function.
2.2.1 Overall Architecture
The PO-CKAN architecture inherits the “Branch-Trunk” structure from DeepONet, as shown in Figure 1. The model has two main components: a branch net and a trunk net. The branch net processes the input function $u(y)$, while the trunk net processes the coordinates $x$. Both nets are built with our novel CKAN layers, which replace traditional MLPs. This substitution is designed to capture complex functional relationships while maintaining computational tractability.
Fig. 1: This figure illustrates the PO-CKAN framework. The model adopts the DeepONet architecture to learn the solution operator mapping input functions to solution functions. The branch and trunk nets are both constructed from our novel CKAN layers. Base Function (Top Panel): The Enhanced Rational Unit (ERU) is used as the computationally efficient and numerically stable base function for all CKAN layers. CKAN Layer (Middle Panel): The CKAN layer is the core innovation. It reduces parameters through a chunk-wise sharing mechanism. Chunks of edges share a single base ERU but each edge retains an individual scalar weight. Training Objective (Bottom Panel): The loss is computed via automatic differentiation and comprises data ($\mathcal{L}_{\text{data}}$), initial condition ($\mathcal{L}_{\text{ic}}$), boundary condition ($\mathcal{L}_{\text{bc}}$), and PDE residual ($\mathcal{L}_{r}$) terms.
2.2.2 CKAN Layer
As previously noted, the KAN architecture’s high expressiveness is coupled with major scalability challenges: the high computational cost of B-spline activations and the prohibitive parameter count. To address these challenges, we introduce the Chunk-rational KAN (CKAN), a novel layer design featuring two key innovations.
Efficient and Stable Rational Activations
To enhance the computational efficiency and expressive power of the KAN layer, we move beyond the traditional B-spline basis and instead employ rational functions. A rational function, which is simply the ratio of two polynomials, offers superior approximation properties for complex behaviors and is highly efficient to compute [27, 10]. Specifically, we parameterize each edge function $\phi(x)$ as a weighted rational function of the form $w\frac{P(x)}{Q(x)}$, where the numerator $P(x)$ and denominator $Q(x)$ are polynomials of degrees $m$ and $n$, respectively.
$\phi(x)=wF(x)=w\cdot\frac{P(x)}{Q(x)}=w\cdot\frac{a_{0}+a_{1}x+\cdots+a_{m}x^{m}}{b_{0}+b_{1}x+\cdots+b_{n}x^{n}}$ (3)
However, the standard form in Eq. 3 is susceptible to numerical instability from poles, which occur where the denominator $Q(x)$ approaches zero. To ensure stable training while retaining high expressiveness, we adopt the Enhanced Rational Unit (ERU) as our basis function, $F(x)$ [51]. The ERU takes the form:
$F(x)=ax+b+\sum^{\frac{n}{2}}_{k=1}\frac{c_{k}x+d_{k}}{(x-e_{k})^{2}+f_{k}^{2}+\epsilon}$ (4)
This formulation, with its always-positive denominator, effectively prevents the occurrence of poles. This choice is further justified by its significant computational advantages. As shown in Table 1, the ERU is substantially more efficient to compute than the original B-spline activations.
Name FLOPs
B-Spline (G=3, K=3) 204
Rational (m=5, n=4) 46
ERU ($n=4$) 19
Table 1: Comparison of FLOPs for different functions. Using the Enhanced Rational Unit reduces FLOPs by approximately 10.7x compared to the B-Spline function.
Chunk KAN
To tackle the challenge of excessive parameters, we introduce a Chunk-wise Parameter Sharing mechanism to create the CKAN. The core concept is to share the base parameters of a rational function across a chunk of connections rather than learning a unique function for each input-output pair. Figure 2 shows the difference between the CKAN, KAN, and a standard MLP. Specifically, for a CKAN layer mapping a $d_{\text{in}}$-dimensional input to a $d_{\text{out}}$-dimensional output:
  1. 1.
    Chunking: We divide the $d_{\text{in}}$ input channels and $d_{\text{out}}$ output channels into $c\times c$ chunks.
  2. 2.
    Parameter Sharing: Within each chunk $(m,n)$, all edges share a single base rational function, $F_{m,n}(x)$, in the form of Eq. 4.
  3. 3.
    Edge-specific Weights: To preserve expressive power, we retain a unique, learnable scalar weight $w_{ij}$ for each individual edge $(i,j)$ in the layer.
Fig. 2: Comparison of our CKAN (with a 2x2 chunk configuration) with a standard KAN and MLP. While a KAN has a unique function on each input-output pair, a CKAN shares a single base function across a chunk of edges.
Suppose $i$ indexes the input channel. With $c\times c$ chunks, the output of the $j$-th neuron in a CKAN layer is given by:
$\text{CKAN}(x)_{j}=\sum_{i=1}^{d_{\text{in}}}w_{ij}\cdot F_{\lfloor i/d_{\text{in},c}\rfloor,\lfloor j/d_{\text{out},c}\rfloor}(x_{i})$ (5)
where $d_{\text{in},c}=d_{\text{in}}/c$ and $d_{\text{out},c}=d_{\text{out}}/c$ are the chunk sizes for the input and output dimensions, respectively, and $(\lfloor i/d_{\text{in},c}\rfloor,\lfloor j/d_{\text{out},c}\rfloor)$ is the 2D index of the chunk to which the edge $(i,j)$ belongs.
In the original KAN formulation, $d_{\text{in}}\times d_{\text{out}}$ distinct activation functions are required. By applying our chunking strategy, this number is reduced to only $c\times c$. In addition to parameter reduction, this approach also decreases computational cost: each input chunk evaluates the base function $F$ once, and the result is reused across the same chunk. Conversely, a vanilla KAN computes a separate $\phi_{i,j}$ for every input–output pair, resulting in substantially higher computational complexity.
Example: Consider a layer with $d_{\text{in}}=d_{\text{out}}=50$, B-spline base function degree $K=5$, and grid size $G=50$. In a vanilla KAN, there are $50\times 50=2500$ unique activation functions. Each base function requires $\text{Func FLOPs}\approx 8K=40$ FLOPs, and the total layer FLOPs is approximately $6.7\times 10^{6}$. For a CKAN with $d_{\text{in}}=d_{\text{out}}=50$, rational degree $4$, and a $2\times 2$ chunk configuration, the total FLOPs is roughly $4.4\times 10^{3}$a reduction by three orders of magnitude.
A comparison of parameter counts and computational requirements is summarized in Table 2.
Name No. Params FLOPs
MLP $d_{in}\times d_{out}+d_{out}$ $\text{Func FLOPs}\times d_{out}+2\times(d_{in}\times d_{out})$
KAN $d_{in}\times d_{out}\times(G+K+3)+d_{out}$ $\text{Func FLOPs}\times d_{in}+(d_{in}\times d_{out})\times[9K(G+1.5K)+2G-2.5K+3]$
CKAN $d_{in}\times d_{out}+d_{out}+(2n+2)\times c^{2}$ $(4.5n+1)\times d_{in}\times c+2\times(d_{in}\times d_{out})$
Table 2: Comparison of parameter counts across different models. Here, Func FLOPs refers to the computational cost of the non-linear activation functions. In KAN, $K$ denotes the order and $G$ the number of grid points. For our CKAN, $n$ indicates the rational order, and $c\times c$ specifies the number of chunks. CKAN maintains a parameter count with only a constant overhead compared to a standard MLP, while the KAN parameters grow with $(G+K+3)$.
2.2.3 Physics-Informed Training Objective
A key challenge in operator learning is that purely data-driven models may yield solutions inconsistent with the underlying physical laws. To address this issue, our PO-CKAN framework embeds the governing physics directly into the training process.
We train the operator network $G_{\theta}$ by minimizing a physics-informed loss function:
$\mathcal{L}(\theta)=\lambda_{\mathrm{data}}\,\mathcal{L}_{\mathrm{data}}(\theta)+\lambda_{\mathrm{ic}}\,\mathcal{L}_{\mathrm{ic}}(\theta)+\lambda_{\mathrm{bc}}\,\mathcal{L}_{\mathrm{bc}}(\theta)+\lambda_{\mathrm{r}}\,\mathcal{L}_{\mathrm{r}}(\theta),$ (6)
where $\lambda_{\ast}\geq 0$ are hyperparameters balancing the different loss components.
Data loss
The data loss $\mathcal{L}_{\mathrm{data}}$ is the Mean Squared Error (MSE) over labeled training pairs:
$\mathcal{L}_{\mathrm{data}}(\theta)=\frac{1}{N_{d}}\sum_{i=1}^{N_{d}}\left\|G_{\theta}(u_{i})(y_{i})-s_{i}(y_{i})\right\|_{2}^{2}.$ (7)
Initial condition loss
The initial condition loss $\mathcal{L}_{\mathrm{ic}}$ enforces the PDE’s initial state $s(y,0)=s_{0}(y)$:
$\mathcal{L}_{\mathrm{ic}}(\theta)=\frac{1}{N_{\mathrm{ic}}}\sum_{j=1}^{N_{\mathrm{ic}}}\left\|G_{\theta}(u_{j})(y_{j}^{\mathrm{ic}},0)-s_{0}(y_{j}^{\mathrm{ic}})\right\|_{2}^{2}.$ (8)
Boundary condition loss
The boundary condition loss $\mathcal{L}_{\mathrm{bc}}$ enforces the PDE’s spatial boundary constraints, here assumed to be Dirichlet for illustration:
$\mathcal{L}_{\mathrm{bc}}(\theta)=\frac{1}{N_{\mathrm{bc}}}\sum_{j=1}^{N_{\mathrm{bc}}}\left\|G_{\theta}(u_{j})(y_{j}^{\mathrm{bc}})-s_{\mathrm{bc}}(y_{j}^{\mathrm{bc}})\right\|_{2}^{2}.$ (9)
PDE residual loss
The physics-informed term $\mathcal{L}_{\mathrm{r}}$ enforces the governing PDE in residual form $\mathcal{R}(u,s)=0$ over collocation points $\{y_{j}^{\mathrm{phys}}\}$:
$\mathcal{L}_{\mathrm{r}}(\theta)=\frac{1}{N_{p}}\sum_{j=1}^{N_{p}}\left\|\mathcal{R}\big(u_{j},G_{\theta}(u_{j})\big)\!\left(y_{j}^{\mathrm{phys}}\right)\right\|_{2}^{2},$ (10)
where all necessary derivatives in $\mathcal{R}$ are obtained via automatic differentiation (AD).
By jointly minimizing these four terms, PO-CKAN learns an operator that is both accurate with respect to observed data and consistent with the initial/boundary conditions and the underlying physical laws.
3 Numerical Experiments
3.1 Burgers’ Equation
We first evaluate our method on the one-dimensional viscous Burgers’ equation [28], a standard benchmark for testing nonlinear wave propagation and shock formation. The governing PDE is given by:
$\begin{cases}\frac{\partial s}{\partial t}+s\,\frac{\partial s}{\partial x}-\nu\,\frac{\partial^{2}s}{\partial x^{2}}=0,&\text{for }x\in(0,1),\ t\in(0,1),\\ s(x,0)=u(x),&\text{for }x\in(0,1),\end{cases}$ (11)
with periodic boundary conditions. The goal is to learn the operator mapping initial conditions $u(x)$ to solutions $s(x,t)$. Initial conditions are sampled from a Gaussian Random Field (GRF), $u\sim\mathcal{N}\left(0,25^{2}(-\Delta+5^{2}I)^{-4}\right)$, to ensure a diverse set of test cases.
[…]
Fig. 3: Modified network architecture
Following the benchmark protocol of [43], we generate 2,000 GRF samples, using 1,500 for training and 500 for testing. Ground truth solutions are obtained via a spectral method implemented in the Chebfun package [16], with 101 temporal snapshots saved per solution. In our experiments, following [54], we adopt the modified network as the base architecture, which incorporates residual connections and gate-controlled mechanisms  3. The modified network exhibits a distinctive forward propagation mechanism [54] as follows
$U$ $=\text{Layer}_{U}(X),\quad V=\text{Layer}_{V}(X),$ (12)
$H^{(1)}$ $=\text{Layer}_{1}(X),$
$Z^{(k)}$ $=\text{Layer}_{k}(H^{(k)}),\quad k=1,\dots,L,$
$H^{(k+1)}$ $=(1-Z^{(k)})\odot U+Z^{(k)}\odot V,\quad k=1,\dots,L,$
$f_{\theta}(x)$ $=\text{Layer}_{\text{out}}(H^{(L)}).$
The PO-CKAN model employs a 4-layer CKAN (rational degree $n=4$, chunk number $c=1x1$) modified network structure for both the branch and trunk networks, with 100 units per layer. For comparison, a baseline PI-DeepONet is configured with the same 4-layer, 100-unit architecture, but using modified MLP layers with $\tanh$ activations. Both models are trained for 100,000 iterations using the Adam optimizer, relying solely on a physics-informed loss that enforces the initial condition, boundary condition, and PDE residuals.
$\mathcal{L}(\theta)=\lambda_{\textbf{ic}}\mathcal{L}_{\text{ic}}(\theta)+\lambda_{\textbf{bc}}\mathcal{L}_{\text{bc}}(\theta)+\lambda_{\textbf{r}}\mathcal{L}_{\text{r}}(\theta),\lambda_{\textbf{ic}}=50,\lambda_{\textbf{bc}}=1,\lambda_{\textbf{r}}=1$ (13)
Experiments are conducted for three viscosity values, $\nu\in\{0.05,0.03,0.01\}$, where smaller $\nu$ leads to sharper solution gradients and more challenging dynamics. The results, presented in Figures 4,5,6,7 and Tables 3-4, indicate a consistent advantage for PO-CKAN. Training loss curves in Figure 4 show faster convergence and lower final loss compared to PI-DeepONet (Table 4). This improved optimization translates into higher predictive accuracy. Qualitative comparisons in Figures 5, 6, and 7 show that PO-CKAN more accurately captures the ground truth, particularly for sharp features at low viscosity. Point-wise error plots confirm this visually, and quantitative results in Table 3 demonstrate substantially lower mean relative $L_{2}$ errors. For $\nu=0.01$, PO-CKAN reduces the error by nearly 48%, highlighting the effectiveness of the CKAN design for learning complex nonlinear operators.
Fig. 4: PO-CKAN vs. PI-DeepONet:loss at different viscosity coefficients $\nu$. ($\nu$ = 0.01, 0.03, 0.05)
Model viscosity $\nu=0.05$ $\nu=0.03$ $\nu=0.01$
PI-DeepONet $1.22\times 10^{-2}$ $1.38\times 10^{-2}$ $6.23\times 10^{-2}$
PO-CKAN $\mathbf{6.93\times 10^{-3}}$ $\mathbf{7.03\times 10^{-3}}$ $\mathbf{3.21\times 10^{-2}}$
Table 3: Mean relative $L^{2}$ errors of PI-DeepONet and PO-CKAN for the Burgers’ equation with different viscosity coefficients $\nu$.
Model viscosity $\nu=0.05$ $\nu=0.03$ $\nu=0.01$
PI-DeepONet $3.41\times 10^{-4}$ $5.59\times 10^{-4}$ $3.66\times 10^{-3}$
PO-CKAN $\mathbf{1.33\times 10^{-4}}$ $\mathbf{1.77\times 10^{-4}}$ $\mathbf{}{1.41\times 10^{-3}}$
Table 4: Final losses of PI-DeepONet and PO-CKAN for the Burgers’ equation with different viscosity coefficients $\nu$.
PI-DeepONet
PO-CKAN
Fig. 5: Comparison of PI-DeepONet and PO-CKAN results for the Burgers’ equation with $\nu=0.05$. The three columns correspond to the ground truth solution (left), the network prediction (middle), and the absolute error (right).
PI-DeepONet
PO-CKAN
Fig. 6: Comparison of PI-DeepONet and PO-CKAN results for the Burgers’ equation with $\nu=0.03$. The three columns correspond to the ground truth solution (left), the network prediction (middle), and the absolute error (right).
PI-DeepONet
PO-CKAN
Fig. 7: Comparison of PI-DeepONet and PO-CKAN results for the Burgers’ equation with $\nu=0.01$. The three columns correspond to the ground truth solution (left), the network prediction (middle), and the absolute error (right).
3.2 Eikonal Equation
We next study a geometric input problem through the two-dimensional Eikonal equation [40], which underlies the computation of Signed Distance Functions (SDFs) widely used for shape representation in computer vision and graphics. Formally, the problem is defined on a domain $\Omega\subset\mathbb{R}^{2}$ as
$\|\nabla s(\mathbf{x})\|_{2}=1,\quad\mathbf{x}\in\Omega,\quad s(\mathbf{x})=0\quad\text{if}\ \mathbf{x}\in\partial\Omega,$ (14)
where $\mathbf{x}=(x,y)$ and the boundary $\partial\Omega$ is a closed curve $\Gamma$. The solution $s(\mathbf{x})$ represents the minimal distance from any point $\mathbf{x}$ in the domain to the boundary $\Gamma$. Our objective is to approximate the solution operator
$G:\Gamma\mapsto s(\mathbf{x}),$
which maps a given boundary shape to its corresponding SDF.
For benchmarking, we consider circles centered at the origin, which allow an analytical solution: the SDF for a circle $\Gamma$ of radius $r$ is given by $s(x,y)=\sqrt{x^{2}+y^{2}}-r$. A training set of $N=1,000$ examples is generated by sampling $r$ uniformly from $U(0.5,1.5)$. Each circle $\Gamma^{(i)}$ is discretized into $m=100$ points, and model predictions are evaluated over the domain $D=[-2,2]\times[-2,2]$.
The PO-CKAN model employs separate 4-layer CKANs (rational degree $n=4$, with a $2\times 2$ chunk configuration) for the branch and trunk networks, each containing 50 units per layer. The baseline PI-DeepONet uses the same 4-layer, 50-unit architecture with standard MLP layers.
Both models are trained for 80,000 iterations using the Adam optimizer with a purely physics-informed loss that enforces the boundary conditions and the Eikonal equation residual condition:
$\mathcal{L}(\theta)=\lambda_{\text{bc}}\mathcal{L}_{\text{bc}}(\theta)+\lambda_{\text{r}}\mathcal{L}_{\text{r}}(\theta),\lambda_{\text{bc}}=\lambda_{\text{r}}=1$ (15)
Here, $\mathcal{L}_{\text{bc}}$ imposes the boundary condition by penalizing deviations of $G_{\theta}(\Gamma^{(i)})(\mathbf{x})$ from zero at boundary points, whereas $\mathcal{L}_{\text{r}}$ penalizes the PDE residual evaluated at $Q=1,000$ collocation points $\mathbf{x}\in\Omega$. In this notation, $\Gamma^{(i)}$ denotes the $i$-th input boundary curve, $\mathbf{x}$ represents a point in the computational domain, and $Q$ is the number of collocation points. The core residual is computed as:
$R_{\theta}^{(i)}(\mathbf{x})=\left(\frac{\partial G_{\theta}(\Gamma^{(i)})}{\partial x}\right)^{2}+\left(\frac{\partial G_{\theta}(\Gamma^{(i)})}{\partial y}\right)^{2}-1.$ (16)
Figure 8 demonstrates a clear performance gap, with PI-DeepONet unable to accurately approximate the operator, exhibiting elevated test loss and reduced prediction accuracy, likely due to its limited network depth. In contrast, PO-CKAN accurately reproduces the circular SDFs, achieving a mean relative $L^{2}$ error of $\mathbf{5.10\times 10^{-3}}$, demonstrating its effectiveness in learning operators from geometric inputs.
Fig. 8: Comprehensive performance comparison between PO-CKAN and PI-DeepONet for learning a circular Signed Distance Function (SDF). Figure (a) shows the test loss curves, where PO-CKAN converges to a loss approximately two orders of magnitude lower than PI-DeepONet ($\sim 10^{-3}$ vs. $\sim 10^{-1}$). Figure (b) displays the results for PI-DeepONet, showing a visually inaccurate prediction and a large absolute error (max 2.5). In contrast, Figure (c) shows the results for PO-CKAN, whose prediction is visually almost identical to the exact solution, with a maximum absolute error of only 0.016.
3.3 Fractional Partial Differential Equations
To further evaluate the model’s capability in handling non-local, history-dependent operators, we consider a time-fractional partial differential equation with a variable-order derivative. Such equations are central to modeling anomalous transport phenomena exhibiting memory effects, including viscoelastic materials and porous media. The governing equation on the domain $t\in[0,1],x\in[0,\pi]$ is:
$\begin{cases}D_{t}^{\alpha(t)}u(x,t)=\frac{\partial^{2}u}{\partial x^{2}}+f(x,t),\\ f(x,t)=\frac{\Gamma(4)}{\Gamma(4-\alpha(t))}t^{3-\alpha(t)}\sin(x)+t^{3}\sin(x),\end{cases}$ (17)
where $D_{t}^{\alpha(t)}$ denotes the variable-order Caputo fractional derivative with linearly varying order $\alpha(t)\in[0,0.9]$, introducing additional temporal complexity[33]. Homogeneous initial ($u(x,0)=0$) and boundary ($u(0,t)=u(\pi,t)=0$) conditions are imposed. To enable precise evaluation, the method of manufactured solutions is employed: $f(x,t)$ is constructed so that the analytical solution is $u(x,t)=\sin(x)t^{3}$, providing a definitive ground truth.
Our PO-CKAN model consists of a branch and a trunk, each implemented as compact 2-layer CKANs (rational degree $n=4$, with a $2\times 2$ chunk configuration) with 20 units per layer, producing 40-unit output vectors. Training is conducted purely in a physics-informed manner without paired input-solution data, guided by a composite loss:
$\mathcal{L}(\theta)=\lambda_{ic}\mathcal{L}_{ic}(\theta)+\lambda_{r}\mathcal{L}_{r}(\theta),\quad\lambda_{ic}=1,\ \lambda_{r}=10,$ (18)
where $\mathcal{L}_{r}$ penalizes the PDE residual,
$R_{\theta}(x,t)=D_{t}^{\alpha(t)}G_{\theta}(u_{0})-\frac{\partial^{2}G_{\theta}(u_{0})}{\partial x^{2}}-f(x,t).$ (19)
The variable-order Caputo derivative is discretized at each time step $t_{k}$ via a convolution formula:
$D_{t}^{\alpha_{k}}u(x,t_{k})\approx\frac{\tau^{-\alpha_{k}}}{\Gamma(2-\alpha_{k})}\sum_{j=1}^{k}c_{j}^{\alpha_{k}}\big[u(x,t_{k-j+1})-u(x,t_{k-j})\big].$ (20)
Figure 9 highlights the model’s strong performance on this challenging fractional problem. Quantitatively, PO-CKAN achieves a mean relative $L^{2}$ error of $\mathbf{2.54\times 10^{-2}}$, representing an improvement of over 80% compared to the baseline ($1.32\times 10^{-1}$). Qualitative comparisons further confirm that our model closely reproduces the reference solution, whereas the baseline shows noticeable deviations. These results demonstrate the architecture’s ability to accurately learn fractional operators and capture complex, history-dependent dynamics.
Fig. 9: Performance comparison of PO-CKAN and PI-DeepONet for solving a fractional PDE (FPDE). The loss curve in (a) shows that while PI-DeepONet’s loss remains high ($\sim 10^{0}$), PO-CKAN’s loss has a low baseline ($\sim 10^{-1}$). The 3D plots show that PI-DeepONet’s prediction remains inaccurate with a maximum error of 0.30 (b). In contrast, PO-CKAN’s prediction of the solution $u(x,t)=\sin(x)\cdot t^{3}$ is visually correct, and its maximum absolute error is exceptionally low at 0.005.
3.4 Diffusion-reaction systems
To conclude our evaluation, we examine a canonical nonlinear problem: a diffusion-reaction system. Such PDEs are fundamental in modeling a variety of physical and biological phenomena, including chemical kinetics and population dynamics. The problem is defined as
$\frac{\partial s}{\partial t}=D\frac{\partial^{2}s}{\partial x^{2}}+ks^{2}+u(x),\quad(x,t)\in(0,1]\times(0,1],$ (21)
subject to homogeneous initial and boundary conditions, with diffusion coefficient $D=0.01$ and reaction rate $k=0.01$.
For this benchmark, PO-CKAN is compared against a baseline PI-DeepONet. Both models employ 5 hidden layers with 50 units each; the key difference lies in the use of CKAN (rational degree $n=4$, with a $2\times 2$ chunk configuration) versus standard MLP layers. Training is fully physics-informed, without requiring paired solution data, and driven by a composite loss:
$\mathcal{L}(\theta)=\lambda_{\text{bc}}\mathcal{L}_{\text{bc}}(\theta)+\lambda_{\text{r}}\mathcal{L}_{\text{r}}(\theta),$ (22)
where $\mathcal{L}_{\text{bc}}$ enforces the zero initial and boundary conditions, and $\mathcal{L}_{\text{r}}$ penalizes the PDE residual,
$R^{(i)}_{\theta}(x,t)=\frac{\partial G_{\theta}(u^{(i)})}{\partial t}-D\frac{\partial^{2}G_{\theta}(u^{(i)})}{\partial x^{2}}-k\big[G_{\theta}(u^{(i)})\big]^{2}-u^{(i)}(x),$ (23)
evaluated at collocation points across the spatio-temporal domain using automatic differentiation.
Figure 10 illustrates the model performance. PO-CKAN exhibits faster convergence and a lower final loss compared to PI-DeepONet. Quantitatively, it achieves a mean relative $L^{2}$ error of $\mathbf{2.58\times 10^{-3}}$, representing an improvement of over 50
Fig. 10: Performance comparison of PO-CKAN and PI-DeepONet on the diffusion-reaction (DR) problem. The loss curves in (a) show that PO-CKAN reaches a final loss ($\sim 10^{-4}$) approximately an order of magnitude lower than PI-DeepONet ($\sim 10^{-3}$). The PI-DeepONet model provides a good visual prediction for the solution $s(x,t)$ (b), with a maximum absolute error of approximately 0.0275. The PO-CKAN model (c) further improves upon this result, yielding a nearly identical visual prediction and reducing the maximum absolute error by a factor of 7, to just 0.004.
3.5 Ablation Studies
We conducted ablation studies to investigate the impact of two key architectural hyperparameters: (i) chunk granularity $c$ and (ii) ERU order $n$. For all experiments, we used identical data splits, optimizers, and training budgets to ensure a fair comparison.
3.5.1 Effect of Chunk Granularity (c)
Table 5 presents the model complexity and inference efficiency for different values of chunk granularity $c$. The results clearly demonstrate that as $c$ increases, both the model’s parameter count and computational load (FLOPs) grow substantially. Specifically, increasing $c$ from 2 to 50 (equivalent to a full KAN) causes the number of parameters to expand by more than 8.5 times (from approximately 23k to 198k) and FLOPs to increase by 7.6 times. In contrast, the impact on inference time is far more moderate, with only a 39% increase from 4.16 ms to 5.79 ms. This highlights a critical trade-off: while a finer chunk granularity (higher $c$) significantly raises the theoretical computational cost, its effect on practical inference latency is modest. This is expected, as higher granularity allows the model to capture more intricate relationships, potentially improving accuracy at the expense of increased model complexity.
Table 5: Model complexity and inference efficiency for different values of $c$. The table reports the number of parameters, FLOPs, and average inference time per sample.
$c$ Params FLOPs Inference (ms)
2 23230 62376 4.164
5 24700 88140 4.392
10 29950 131080 4.612
25 66700 259900 5.140
50 (full KAN) 197950 474600 5.789
3.5.2 Effect of ERU Order (n)
Table 6 investigates the impact of the ERU order $n$ on model accuracy, with a fixed chunk granularity of $c=2$. The results show a clear trend: increasing $n$ leads to a consistent and significant reduction in the relative $L^{2}$ error. For instance, by raising $n$ from 4 to 20, the error decreases substantially from $6.23\times 10^{-2}$ to $1.19\times 10^{-2}$, an improvement of over 80%. Notably, this substantial gain in accuracy comes at a negligible cost in model complexity. Over the same range of $n$, the parameter count increases minimally from 23,230 to 24,126, a rise of less than 4%. This experiment powerfully demonstrates that increasing the ERU order is a highly parameter-efficient strategy for enhancing model performance, allowing for a major boost in function approximation capability with a minimal increase in model size.
Table 6: Effect of ERU order $n$ with fixed $c=2$. The final column indicates the mean relative $L^{2}$ error.
$n$ Params Relative $L^{2}$ Error
4 23230 $6.23\times 10^{-2}$
8 23454 $2.33\times 10^{-2}$
12 23678 $2.26\times 10^{-2}$
16 23902 $1.27\times 10^{-2}$
20 24126 $1.19\times 10^{-2}$
4 Conclusion and Future Work
In this work, we introduced the Physics-Informed Deep Operator CKAN (PO-CKAN), a novel framework for learning physically consistent solution operators for partial differential equations. By combining the physics-informed training paradigm with a new, highly efficient Chunk-rational KAN (CKAN) architecture, our model demonstrates superior accuracy and computational efficiency compared to standard physics-informed operator networks. The CKAN layer, with its stable rational activations and chunk-wise parameter sharing, proved effective at capturing complex, nonlinear operator dynamics across a range of benchmark problems.
Building on this foundation, several promising research directions emerge. First, an adaptive version of the CKAN layer could be developed, where the chunk granularity or the order of the rational functions is automatically refined based on the complexity of the learned operator. This could further enhance both accuracy and efficiency, particularly for problems with sharp, localized features or multiscale phenomena. Second, extending the PO-CKAN framework to problems with more complex geometries and boundary conditions, for instance by integrating it with domain decomposition methods or mesh-free collocation strategies, would significantly broaden its applicability to real-world engineering simulations [4].
A third direction involves the integration of uncertainty quantification (UQ) methodologies [55]. Developing a Bayesian formulation of PO-CKAN would enable the model to handle noisy or sparse observational data and provide robust confidence intervals for its predictions, which is critical for safety-conscious applications. Finally, leveraging the inherent parallelism of the chunk-based architecture for high-performance computing (HPC) represents a crucial next step. Optimizing the CKAN layer for distributed GPU or TPU execution could unlock its potential for large-scale, high-fidelity simulations, paving the way for real-time PDE solvers in demanding fields like climate science and computational fluid dynamics.
Acknowledgment
This work was supported by the National Science Foundation (NSF) under grants DMS-2053746, DMS-2134209, ECCS-2328241, CBET-2347401 and OAC-2311848, and by the U.S. Department of Energy (DOE) Office of Science Advanced Scientific Computing Research program under award number DE-SC0023161, and the DOE–Fusion Energy Science program, under grant number: DE-SC0024583.